### v1

# Designing Informative Metrics for Few-Shot Example Selection

Abstract
Pretrained language models (PLMs) have shown remarkable few-shot learning capabilities when provided with properly formatted examples. However, selecting the “best” examples remains an open challenge. We propose a complexity-based prompt selection approach for sequence tagging tasks. This approach avoids the training of a dedicated model for selection of examples, and instead uses certain metrics to align the syntactico-semantic complexity of test sentences and examples. We use both sentence- and word-level metrics to match the complexity of examples to the (test) sentence being considered. Our results demonstrate that our approach extracts greater performance from PLMs: it achieves state-of-the-art performance on few-shot NER, achieving a 5% absolute improvement in F1 score on the CoNLL2003 dataset for GPT-4. We also see large gains of upto 28.85 points (F1/Acc.) in smaller models like GPT-j-6B. [^1] Structured prompting base code taken from https://github.com/blvns/structured_prompting.git.
Designing Informative Metrics for Few-Shot Example Selection
Rishabh Adiga University of Illinois Urbana-Champaign radiga2@illinois.edu                        Lakshminarayanan Subramanian New York University lakshmi@nyu.edu
Varun Chandrasekaran University of Illinois Urbana-Champaign varunc@illinois.edu
1 Introduction
Pretrained language models (PLMs) have demonstrated impressive few-shot learning capabilities when prompted with examples for a particular task (Brown et al., 2020). In few-shot prompting, $k$ examples that demonstrate the goal of the downstream task along with the test sample (collectively called the $k$-shot prompt) are provided to the PLM. But the effectiveness of PLMs in few-shot learning depends on the quality of these prompts; this is challenging due to the need for unambiguous instruction (Sclar et al., 2024) which is further complicated due to the task of selecting examples. Appropriately selected examples will ensure that the PLM can generalize effectively and perform optimally across diverse natural language processing (NLP) tasks.
Recent work has shown that specific examples provided in-context can significantly impact the performance of PLMs (Min et al., 2022)Sorensen et al. (2022) notes that specific example combinations lead to better generalization, and connects this with the informativeness of these examples. This suggests that surfacing specific properties of prompts result in improved performance. Most prior work for prompt selection has either trained a prompt retriever to select examples, or used PLMs themselves to label examples (Rubin et al., 2022) to serve as a good prompt. However, both approaches require additional computation and task specific training. Liu et al. (2022) uses nearest neighbours algorithms to select examples. However, this approach solely relies on embedding space properties, and does not consider other metrics related to the label distribution, leading to sub-optimal performance in some tasks.
In this work, we propose a new approach for example selection for sequence tagging tasks. These tasks are important because they enable the extraction of structured information from unstructured text data; this serves as a fundamental building block for various NLP tasks. Our approach, complexity-based prompt retrieval (CP retrieval), aims to select examples from the training data whose “properties” align with that of the test sentence. To this end, we use sentence- and word-level complexity measures to match prompt examples to evaluation sentences.
We conduct our experiments on named entity recognition (NER), part-of-speech (POS) tagging, and sentence chunking. We also devise a new dataset for privacy policies involving tagging of contextual integrity parameters (Shvartzshnaider et al., 2019). We show that CP retrieval improves accuracy significantly in all cases.
The contributions of this work are as follows:
  1. 1.
    We propose CP retrieval to select prompts based on sentence and word level complexity metrics (§ 3).
  2. 2.
    We demonstrate improved few-shot accuracy with CP retrieval. Our approach achieves state-of-the-art performance in few-shot NER with a 5% absolute improvement in F1 score on the CoNLL2003 dataset for GPT-4 and upto 28.85 points (F1/Accuracy) in smaller models (§ 5).
2 Related Work
To perform our experiments, we use the Structured Prompting paradigm introduced by Blevins et al. (2023) as our baseline for named entity recognition (NER), part-of-speech (POS) tagging, and sentence chunking. Each model that we use is provided $k$ pairs of (sentence, tagged sentence) as the task examples in a $k$-shot prompt. This involves iteratively feeding the PLM a word and its predicted label to get it to label the next word. These are sequence tagging tasks—those that assign labels or tags to tokens or other units of text. Experiments show that PLMs can effectively solve these tasks with just a few demonstrations using this technique.
Liu et al. (2022) has explored strategies for selecting examples to leverage GPT-3’s few-shot learning capabilities. They analyzed the sensitivity of GPT-3’s performance to randomly sampled examples and found significant variability (a phenomenon also explained by Min et al. (2022)).  Liu et al. (2022) proposed a retrieval based approach called KATE that selects semantically similar examples by using $k$-nearest neighbors in the embedding space of pre-trained sentence encoders like BERT and RoBERTa. They found that KATE substantially improved GPT-3’s results over random sampling on natural language understanding and generation tasks. Results with the $k$NN based approach as baseline are demonstrated in Appendix A.2.
3 Methodology
Our core idea is simple: to select examples that closely resemble the test sample in both semantics and length while ensuring a diverse range of labels for the task. We measure the “complexity” of a candidate example using 3 different sentence-level and word-level metrics: (a) normalized sentence similarity score, (b) normalized smoothed length similarity, and (c) normalized label entropy.
Problem Setup: Assume we have a training dataset of the form $D=\{z_{i}=(x_{i},y_{i})\}_{i=1}^{n}$ and a test example $x_{test}$, our goal is to use a complexity score to select a subset of examples $S$ of the training dataset ($S\subset D$) s.t. $|S|=k$.
3.1 Normalized Sentence Similarity Score
For a given test sentence, we calculate its similarity scores with each candidate sentence for the $k$-shot prompt. We convert each of the sentences into sentence embeddings using the all-MiniLM-L6-v2 sentence transformer (results do not vary drastically with alternate embedding methods). This produces an embedding size of 384 dimensions. To find the similarity score, we calculate cosine similarity between the embeddings.
$\texttt{Similarity}(x_{i},x_{test})=\texttt{cosine\_sim}(\texttt{emb}(x_{i}),\texttt{emb}(x_{test}))$
These values are normalized to get our final sentence similarity scores. Formally, this is denoted as NormSentSimilarity [^2] Normalization here is N($s_{i}$) = $s_{i}$ / max($S$) where $S=\{s_{1},\cdots,s_{n}\}$ and $s_{i}\in\mathbb{R}$.
3.2 Normalized Smoothed Length Similarity
This metric is based on the difference in sentence lengths between the candidate sentence and the test sentence. We use this metric since cosine similarity of sentence embeddings does not inherently factor sentence length. The smoothed length similarity (SLS) transforms the absolute length difference between two sentences using a sigmoid function to produce a smooth, tapered similarity curve rather than a hard threshold. The sigmoidal nature of SLS prevents stark drops in scores for small length divergences. This is important as, for few-shot prompting, some flexibility in length ranges is desirable to provide syntactic variety while maintaining coarse length matching. It is formulated as follows:
$\texttt{SLS}(x_{i},x_{test})=(1+\exp({\dfrac{|\texttt{len}(x_{i})-\texttt{len}(x_{test})|}{T}}))^{-1}$
where $T$ controls the sigmoidal shape ($T$ = 3 in our experiments). Sentences with similar lengths achieve higher SLS scores, while increasingly divergent lengths taper down towards 0 in a continuous, non-binary fashion. These values are then normalized to provide our final Normalized Smoothed Length Similarity scores denoted as NormSLS22footnotemark: 2.
3.3 Normalized Label Entropy
For a given task, let $\mathcal{Y}$ define the space of possible labels. For sentence $x_{i}$, let $y_{i}$ be the list of labels, and $\nu({y}^{j}_{i})$ is the frequency of label $j\in\mathcal{Y}$ in list ${y}_{i}$. We can calculate $\Pr({y}^{j}_{i})=\nu({y}^{j}_{i})/\texttt{len}(x_{i})$. Entropy is then calculated as:
$\text{H}({y}_{i})=-\sum_{j\in\mathcal{Y}}\Pr({y}^{j}_{i})\log_{2}\Pr({y}^{j}_{i})$
The intuition is that sentences where labels are more skewed provide less information gain for in-context learning compared to flatter distributions. Entropy can be used to quantify that skew difference. We normalize these entropy values to get our final NormEntropy22footnotemark: 2 scores.
3.4 Complexity Score
We propose a complexity score to align the syntactico-semantic complexity of prompts and examples. The three component metrics are weighted and summed to produce the final complexity score:
$\texttt{CS}(z_{i},x_{test})=w_{1}*\texttt{NormSLS}(x_{i},x_{test})$
$+w_{2}*\texttt{NormEntropy}(y_{i})$
$+w_{3}*\texttt{NormSentSimilarity}(x_{i},x_{test})$
We use this score to select the $k$ highest scoring train sentences for a $k$-shot prompt. $w_{1}$, $w_{2}$, $w_{3}$ are weights that are set using grid search to optimize scores on the development set for each sequence tagging task. We found that $(w_{1},w_{2},w_{3})$ = $(0.25,0.25,0.5)$ is the best value for NER, $(w_{1},w_{2},w_{3})$ = $(0.2,0.1,0.7)$ is the best value for Chunking and $(w_{1},w_{2},w_{3})$ = $(0.1,0.1,0.8)$ is the best value for POS Tagging. We see that the above set of values indicate that sentence similarity is more important. This is most likely due to the presence of extremely similar examples in the test and training set which are flaws with respect to the dataset.
4 Experimental Setup
4.1 Tasks
We perform our experiments on 4 different sequence tagging tasks namely NER, POS tagging, sentence chunking and on a custom dataset of privacy policies for tagging contextual integrity (CI) parameters (Shvartzshanider et al., 2023). Each task involves the annotation of tokens in a sequence with specific labels, facilitating the extraction of valuable information and enhancing language understanding. For all datasets, we use a random set of 1000 test samples. For the CI task, we use 89 samples.
1. NER: This focuses on classifying named entities within a text. The entities can range from persons and organization to locations and dates. For our experiments, we use the widely recognized CoNLL2003 dataset (Tjong Kim Sang and De Meulder, 2003) for measuring the proficiency of our CP retrieval.
2. POS Tagging: This involves assigning grammatical categories (e.g., nouns, verbs, adjectives) to each word in a sentence. We evaluate the POS tagging performance using the English Universal Dependencies (UD) treebank annotated on the GUM corpus (Zeldes, 2017). It employs the UPOS tagset introduced by Nivre et al. (2020).
3. Sentence Chunking: This aims to partition sentences into non-overlapping syntactic units. For our experiments, we use the CoNLL2000 dataset (Tjong Kim Sang and Buchholz, 2000) to frame chunking as a BIO tagging task.
4. CI parameters: We additionally created a custom dataset (which we will release) that involves the sequence tagging task of CI parameters for privacy policies which is taken from Shvartzshanider et al. (2023). CI parameters sequence tagging is a crucial task within the realm of privacy-aware NLP (Shvartzshnaider et al., 2019).
4.2 Models
We perform our experiments on the GPT-Neo series of models which has parameter counts from 125 million to 20 billion (GPT-Neo-125M,GPT-Neo-1.3B, GPT-Neo-2.7B, GPT-j-6B, GPT-NeoX-20B) and additionally 2 black box models (Davinci-002 and GPT-4) for the sequence tagging tasks of NER, POS tagging and sentence chunking. For our custom dataset for CI parameters tagging, we use Llama2 variants.
NER (F1) Chunking (F1) POS (Acc.)
Baseline CP retrieval Baseline CP retrieval Baseline CP retrieval
GPT-Neo-125M $12.46$ 29.21 $24.28$ 40.33 $55.53$ 71.09
GPT-Neo-1.3B $31.49$ 52.44 $25.44$ 44.19 $65.30$ 75.58
GPT-Neo-2.7B $25.73$ 41.77 $28.50$ 53.32 $64.68$ 76.19
GPT-j-6B $25.88$ 54.73 $35.85$ 54.56 $79.96$ 84.37
GPT-Neox-20B $37.74$ 61.49 $56.18$ 59.66 $80.64$ 86.21
Davinci-002 (Black Box) $20.30$ 40.03 $35.82$ 51.79 46.0446.0446.0446.04333Accuracy has changed with respect to the original paper for this model probably because of the the reasoning in Chen et al. (2023) 52.20
GPT-4 (Black Box) $83.48$* 88.76 $79.78$ 84.52 $92.93$ 93.47
Table 1: Results of using complexity score retrieval (CP retrieval) of the best scoring $k$ examples for the few-shot prompt (here $k$ = 5). The baseline is in the structured prompting paradigm (Blevins et al., 2023) with static prompt examples. CP retrieval significantly surpasses the baseline using linguistic structure alone for the prompt (details in Appendix A.1) in all cases. Best values are in boldface. * in the table shows the previously achieved state-of-the-art performance on few-shot NER by Ashok and Lipton (2023).
CI tagging (Acc.)
Baseline CP retrieval
Llama2-13B $37.56$ 56.79
Llama2-70B $43.90$ 59.35
Table 2: Results of using complexity score for retrieval (CP retrieval) on the CI parameters tagging task using Llama2 models. Best values are in boldface.
Figure 1: CP retrieval demonstrates superior performance across all tasks. (a) Performance while varying number of demonstrations $k$ (GPT-Neo-2.7B) for NER. (b) Performance while varying model size for a fixed $k$ = 5 for POS. (c) Performance for the various tasks on GPT-NeoX-20B model with $k$ = 5. Note taht NER and chunking are evaluated based on the F1 score, and POS is evaluated based on accuracy as commonly done.
5 Results
Our results show that CP retrieval significantly improves few-shot accuracy over the baseline across all models and considered tasks. Table 1 compares the F1 and Accuracy scores (for NER, chunking and POS tagging) for different models while keeping the number of examples constant. Table 2 provides the same for the CI tagging task on Llama models.
We observe substantial gains using this method with the largest gains on GPT-j-6B which achieves a 28.85 point increase on NER, GPT-Neo-2.7B which achieves a 24.82 point increase in chunking and GPT-Neo-125M which achieves a 15.56 point increase in POS tagging. It is also evident that the gains of example selection drops with increasing model size in figures 1(b) and 1(c) (with GPT-4 having the least gains) and this is due to the method approaching overall accuracy saturation points. This is also combined with the fact that larger models have seen more diverse training data, so random prompt selections are more than likely to cover the needed distribution and hence, additional gains from optimal selection are diminished. However, even with small gains, we achieve state-of-the-art performance on the CoNLL2003 dataset with a 5% increase over PromptNER methodology (Ashok and Lipton, 2023) for few-shot NER .
From Figure 1(a), we see that our approach enables effective learning (larger gains) with smaller $k$. At $k$ = 2, we see a gain of 19.37 points over the baseline which decreased when $k$ is increased. The percentage gains consistently reduce as we increase the number of examples.
Discussion: While we have focused on complexity metrics for selecting examples in simple prompting setups, what we want to put forth is that this approach can further boost performance when combined with more sophisticated prompting techniques. For example, CP retrieval can be used along with chain-of-thought (CoT) prompting (Wei et al., 2023). It can be used to identify the most information-rich examples for the initial context in CoT prompting which will result in larger gains.
6 Conclusions
In summary, CP retrieval is a flexible technique that can be used to enhance the accuracy of existing few-shot approaches using example scoring. Additionally, it provides a task-agnostic quantification of the informativeness of examples for improved model performance. The relative weights of the components of the complexity score can be tuned on a per-task basis. Our approach allows us to extract the most out PLMs without any fine-tuning of model parameters.
7 Limitations
Our methodology focuses on sequence tagging tasks and cannot be applied to tasks outside this domain such as question answering. This is because the normalized entropy metric is specifically designed for encouraging selection of sentences with a wider label variety on the token level; this helps increase the sampling of underrepresented labels. Further, we have conducted all experiments for the sequence tagging tasks in only English language. Performance with other languages has not been analyzed.
Appendix A Appendix
A.1 Example Prompts
Context: The Finance Ministry raised the price for tap sales of the Dutch government ’s new 5.75 percent bond due September 2002 to 99.95 from 99.90 . Tagged: The_O Finance_B-ORG Ministry_I-ORG raised_O the_O price_O for_O tap_O sales_O of_O the_O Dutch_B-MISC government_O ’s_O new_O 5.75_O percent_O bond_O due_O September_O 2002_O to_O 99.95_O from_O 99.90_O ._O Context: Swiss bonds ended mostly higher in generally quiet activity , with the September confederate bond futures contract holding just above 113.00 . Tagged: Swiss_B-MISC bonds_O ended_O mostly_O higher_O in_O generally_O quiet_O activity_O ,_O with_O the_O September_O confederate_O bond_O futures_O contract_O holding_O just_O above_O 113.00_O ._O Context: The Brent crude futures market on the Singapore International Monetary Exchange ( SIMEX ) was closed on Monday in respect for a U.K. national holiday . Tagged: The_O Brent_B-ORG crude_O futures_O market_O on_O the_O Singapore_B-ORG International_I-ORG Monetary_I-ORG Exchange_I-ORG (_O SIMEX_B-ORG )_O was_O closed_O on_O Monday_O in_O respect_O for_O a_O U.K._B-LOC national_O holiday_O ._O Context: European bourses closed mixed on Tuesday with London clawing back most of the day ’s losses despite an unsteady start on wall Street , hit by inflation worries . Tagged: European_B-MISC bourses_O closed_O mixed_O on_O Tuesday_O with_O London_B-LOC clawing_O back_O most_O of_O the_O day_O ’s_O losses_O despite_O an_O unsteady_O start_O on_O wall_B-ORG Street_I-ORG ,_O hit_O by_O inflation_O worries_O ._O Context: No closures of airports in the Commonwealth of Independent States are expected on August 24 and August 25 , the Russian Weather Service said on Friday . Tagged: No_O closures_O of_O airports_O in_O the_O Commonwealth_B-LOC of_I-LOC Independent_I-LOC States_I-LOC are_O expected_O on_O August_O 24_O and_O August_O 25_O ,_O the_O Russian_B-ORG Weather_I-ORG Service_I-ORG said_O on_O Friday_O ._O Context: Hungarian overnight interest rates closed higher on Friday as market liquidity tightened before the December 10 social security contribution payment deadline, dealers said. Tagged:
Figure 2: Prompt generated for a specific test sample after retrieval of 5 examples (5-shot prompt) from the training dataset for Named Entity Recognition using our approach. The test sentence that we want the model to label is provided at the end of the prompt and is shown in bold
Context: Shearson is offering the notes as 6 34 % securities priced to yield 6.15 % . Tagged: Shearson_B-NP is_B-VP offering_I-VP the_B-NP notes_I-NP as_B-PP 6_B-NP 34_I-NP %_I-NP securities_I-NP priced_B-VP to_B-VP yield_I-VP 6.15_B-NP %_I-NP ._O Context: Bonds due 1991-1996 carry 6.70 % coupons and bonds due 1997-2000 carry 6 34 % coupons . Tagged: Bonds_B-NP due_B-ADJP 1991-1996_B-NP carry_B-VP 6.70_B-NP %_I-NP coupons_I-NP and_O bonds_B-NP due_B-ADJP 1997-2000_B-NP carry_B-VP 6_B-NP 34_I-NP %_I-NP coupons_I-NP ._O Context: Their price falls further than that of other bonds when inflation and interest rates kick up . Tagged: Their_B-NP price_I-NP falls_B-VP further_B-ADVP than_B-PP that_B-NP of_B-PP other_B-NP bonds_I-NP when_B-ADVP inflation_B-NP and_O interest_B-NP rates_I-NP kick_B-VP up_B-PRT ._O Context: Serial bonds were priced at par to yield to 6.90 % in 2000 . Tagged: Serial_B-NP bonds_I-NP were_B-VP priced_I-VP at_B-PP par_B-NP to_B-VP yield_I-VP to_B-PP 6.90_B-NP %_I-NP in_B-PP 2000_B-NP ._O Context: At the auction of six-month U.S. Treasury bills on Monday , the average yield fell to 7.61 % from 7.82 % . Tagged: At_B-PP the_B-NP auction_I-NP of_B-PP six-month_B-NP U.S._I-NP Treasury_I-NP bills_I-NP on_B-PP Monday_B-NP ,_O the_B-NP average_I-NP yield_I-NP fell_B-VP to_B-PP 7.61_B-NP %_I-NP from_B-PP 7.82_B-NP %_I-NP ._O Context: The rate on six-month bills rose to 7.53 % for a bond-equivalent yield of 7.92 % . Tagged:
Figure 3: Prompt generated for a specific test sample after retrieval of 5 examples (5-shot prompt) from the training dataset for Sentence Chunking using our approach. The test sentence that we want the model to label is provided at the end of the prompt and is shown in bold
Context: Many forms of culture are passed down through a combination of deliberate and unconscious processes . Tagged: Many_ADJ forms_NOUN of_ADP culture_NOUN are_AUX passed_VERB down_ADP through_ADP a_DET combination_NOUN of_ADP deliberate_ADJ and_CCONJ unconscious_ADJ processes_NOUN ._PUNCT Context: Cognitive psychology is the field of psychology dedicated to examining how people think . Tagged: Cognitive_ADJ psychology_NOUN is_AUX the_DET field_NOUN of_ADP psychology_NOUN dedicated_VERB to_SCONJ examining_VERB how_ADV people_NOUN think_VERB ._PUNCT Context: Having been a psychologist for a number of years gives me a leg up on it . Tagged: Having_AUX been_AUX a_DET psychologist_NOUN for_ADP a_DET number_NOUN of_ADP years_NOUN gives_VERB me_PRON a_DET leg_NOUN up_ADV on_ADP it_PRON ._PUNCT Context: Do they get mad or irritated if the centre of attention moves to someone else ? Tagged: Do_AUX they_PRON get_VERB mad_ADJ or_CCONJ irritated_VERB if_SCONJ the_DET centre_NOUN of_ADP attention_NOUN moves_VERB to_ADP someone_PRON else_ADV ?_PUNCT Context: When thoughts are formed , the brain also pulls information from emotions and memories ( Figure 7.2 ) . Tagged: When_ADV thoughts_NOUN are_AUX formed_VERB ,_PUNCT the_DET brain_NOUN also_ADV pulls_VERB information_NOUN from_ADP emotions_NOUN and_CCONJ memories_NOUN (_PUNCT Figure_PROPN 7.2_NUM )_PUNCT ._PUNCT Context: Psychologists have examined the mental processes that underpin conscious and unconscious biases [ 6 ] ; Tagged:
Figure 4: Prompt generated for a specific test sample after retrieval of 5 examples (5-shot prompt) from the training dataset for Part-of-Speech tagging using our approach. The test sentence that we want the model to label is provided at the end of the prompt and is shown in bold
A.2 $k$NN Baseline
NER (F1) Chunking (F1) POS (Acc.)
$k$NN Baseline CP retrieval $k$NN Baseline CP retrieval $k$NN Baseline CP retrieval
GPT-Neo-125M $25.39$ 29.21 $38.16$ 40.33 $66.82$ 71.09
GPT-Neo-2.7B $36.44$ 41.77 $49.05$ 53.32 $73.71$ 76.19
Table 3: Results of using complexity score retrieval (CP retrieval) of the best scoring $k$ examples for the few-shot prompt (here $k$ = 5). The baseline here is picking the $k$-nearest neighbors + structured prompting paradigm (Blevins et al., 2023).